QUESTION 3

  1. Clearly explain the features of Fixed Income Securities. Illustrate your response with a relevant example from the Zambian financial markets.

                                                                                            [10 Marks]

  1. A 364 Days Treasury bill of K1,000 Face Value is currently on offer at K840.33.

  1. Calculate its Yield to maturity (YTM) at this price.

                                                                        [03 Marks]

  1. Assuming the market yields for similar Bills were 200 basis points higher and lower, calculate the trading price of the bill under each case.

                                                                                 [03 Marks]

  1. Briefly comment on the relationship between the price of the Bill and its YTM exhibited above.

                                                                                 [04 Marks]

Total 20 Marks
Answer #1

a]

Fixed income securities have the following features :

  • Fixed income securities have a maturity date, when the principal is repaid
  • Par value - they have a par value, or face value, which represents the principal borrowed
  • They may be coupon securities, in which case they pay period coupon payments, or they may be zero-coupon securities, in which case they are issued at a discount to par value and there are no periodic coupon payments

For example, a Zambia Government bond maturing in April 2024 carries a coupon rate of 8.5% and a par value of $1,000.

International bonds: Zambia 8.5% 14apr2024 USD (XS1056386714, 988895AE8) Status Country of risk Maturity (option) Amount o ou

b]

i]

Price of bill = par value / (1 + YTM)years to maturity

840.33 = 1000 / (1 + YTM)1

YTM = (1000 / 840.33) - 1

YTM = 19.00%

ii]

If yields are 200 basis points higher

YTM = 19.00% + 2% = 21%

Price of bill = par value / (1 + YTM)years to maturity

Price of bill = 1000 / (1 + 21%)1

Price of bill = K826.45

If yields are 200 basis points lower

YTM = 19.00% - 2% = 17%

Price of bill = par value / (1 + YTM)years to maturity

Price of bill = 1000 / (1 + 17%)1

Price of bill = K854.70

iii]

As exhibited above, the YTM and bill price have an inverse relationship.

Higher the YTM, lower the bill price and lower the YTM, higher the bill price.